### v1

# IBIS: A Powerful Hybrid Architecture for Human Activity Recognition

Abstract
The increasing interest in Wi-Fi sensing stems from its potential to capture environmental data in a low-cost, non-intrusive way, making it ideal for applications like healthcare, space occupancy analysis, and gesture-based IoT control. However, a major limitation in this field is the common problem of overfitting, where models perform well on training data but fail to generalize to new data. To overcome this, we introduce a novel hybrid architecture that integrates Inception-BiLSTM with a Support Vector Machine (SVM), which we refer to as IBIS. Our IBIS approach is uniquely engineered to improve model generalization and create more robust classification boundaries. By applying this method to Doppler-derived data, we achieve a movement recognition accuracy of nearly 99%. Comprehensive performance metrics and confusion matrices confirm the significant effectiveness of our proposed solution.
I Introduction
Wi-Fi sensing is an increasingly important technology that leverages ubiquitous Wi-Fi signals for non-intrusive human activity Recognition (HAR) [1]. Unlike traditional methods that rely on cameras or wearable sensors, Wi-Fi sensing utilizes Channel State Information (CSI) to capture and analyze subtle changes in signal propagation caused by human movement [2]. This approach is not only cost-effective but also preserves user privacy, making it highly suitable for applications in healthcare, surveillance, and smart homes.
However, a key challenge in this field is overfitting, where models trained on a specific dataset fail to generalize to new environments or unseen data. Many existing solutions, particularly those based on Convolutional Neural Networks (CNNs) [3] [4], are effective at capturing local patterns but often struggle to learn the broader, temporal dependencies critical for robust generalization. Furthermore, raw CSI data frequently contains noise from hardware artifacts, such as phase and frequency oscillations, which can significantly hinder accurate movement interpretation and requires a rigorous sanitization process.
To address these limitations, we propose a novel hybrid neural network architecture named Inception-BiLSTM (IBIS). The Inception component efficiently extracts local features from the input data [5], while the Bidirectional Long Short-Term Memory (BiLSTM) component captures long-range temporal dependencies. To further enhance the model’s predictive power and refine classification boundaries, we integrate a SVM with a non-linear kernel in a post-processing stage. This ensemble learning approach [6] leverages the strengths of multiple algorithms to achieve a more robust and reliable solution.
Our model’s effectiveness was evaluated on five and eight human activity classification tasks using Doppler-derived data. We compared our IBIS approach against two state-of-the-art methods: the SHARP algorithm [7], a recent ensemble-based method for HAR, and an Attention-Based BiLSTM (ABLSTM) approach [8] that uses raw CSI. The results demonstrate that our hybrid Inception-BiLSTM network, augmented with a non-linear SVM, significantly outperforms both benchmark models. Our method achieved an average accuracy of nearly 99%, validating the benefits of combining Doppler features with a sophisticated hybrid deep learning architecture for enhanced HAR performance, particularly in challenging generalization scenarios.
The remainder of this paper is structured as follows. Section II presents a comprehensive overview of existing research on HAR using hybrid neural networks, providing a fresh perspective on the field. Section III details the development and implementation of our proposed algorithms, including the specific neural network architecture. The validation process, which includes activity classification confusion matrices and Receiver Operating Characteristic (ROC) curve analysis, is described in Section IV. Finally, Section V provides a summary of our findings and offers concluding remarks.
II Related Works
This section provides a review of state-of-the-art research in Wi-Fi sensing, focusing on the machine learning algorithms commonly used for CSI data analysis. We also discuss ensemble learning techniques that have been applied to optimize model performance. Recent studies have explored various approaches to enhance Wi-Fi sensing capabilities. Table I presents a comparative overview of the models and key features from previous studies on CSI-based Human Activity Recognition.
Son et al. [9] improved indoor occupancy detection by proposing a preprocessing method that uses nonlinear transformations to expand CSI data into three channel vectors, which amplified signal differences. A CNN was then applied, achieving an average accuracy of approximately 95%. This work highlights the importance of effective signal preprocessing for boosting classification performance.
Ensemble learning has emerged as a key strategy to improve the reliability and accuracy of machine learning models. Wei et al. [10] used an ensemble learning approach for HAR by combining feature extraction with an SVM classifier. This method reportedly outperformed other literature approaches, reaching an accuracy of about 98.95%.
Similarly, Ye et al. [11] developed a system for intrusion monitoring that combines Deep Neural Networks (DNNs) with BiLSTM networks. While their approach demonstrated improved performance over traditional methods, they noted a significant weakness in detecting specific actions, such as falling movements, with an accuracy of only 62.65%.
The SHARP algorithm, as proposed by Meneghello et al. [7], provides a foundational method for HAR based on Doppler traces. Their work addressed noise and hardware interference and, after applying a neural network-based refinement, achieved an accuracy of up to 95%. The authors also demonstrated that antenna merging could further enhance accuracy. Building upon this, Cominelli et al. [12] investigated the impact of Wi-Fi 6 features and developed a large-scale dataset for benchmarking. However, they reported difficulties in reproducing the SHARP algorithm’s results, suggesting a sensitivity to environmental variability. Their findings also indicated that SHARP’s performance degrades with increased bandwidth and a higher number of activity classes.
Despite these advancements, a major challenge in Wi-Fi sensing remains the issue of poor generalization to unseen data. He et al. [8] addressed this limitation by introducing a cross-subject transfer learning method based on a CNN-ABLSTM model. By training on separate source and target domains, they achieved an accuracy of 85.82% in cross-subject recognition tasks, demonstrating a promising path for creating more robust and generalizable models.
TABLE I: Comparison of Models and Features in Human Activity Recognition Using CSI
Specification [9] [10] [7] [12] [11] [8]
Model CNN CNN+RF+SVM Inception Inception DNN-ABLSTM CNN-ABLSTM
Accuracy 95% 98.85% 88.12% 77.72% 92.06% 85.82%
Application Indoor Occupancy HAR HAR HAR HAR HAR
Network 802.11 802.11n 802.11ac 802.11ac/ax 802.11n -
Bandwidth 20MHz - 80MHz 80/160MHz - -
CSI Data Amplitude Temporal and Spatial Correlation Phase CFR Phase CFR Amplitude and Phase Raw CSI
III System Modeling
III-A Dataset
The dataset for this research is based on the SHARP study [7] and consists of CSI data collected from three volunteers (one male, two females) performing eight activities in various indoor environments. These activities: empty, sitting, stand up, walking, running, jumping, and arm gym—were each recorded for 120 seconds. The data was gathered in a bedroom, living room, and university laboratory during two periods: April–December 2020 and January 2022. A Nexmon CSI Extraction Tool and an Asus RT-AC86U Wi-Fi router were used to extract the CSI from Wi-Fi 2.4 GHz preamble headers.
CSI is a data framework that describes how a wireless signal travels from a transmitter to a receiver. It provides high-resolution data on the physical propagation properties of a signal, including the effects of diffraction, reflection, and scattering [9]. Because CSI captures both amplitude and phase information, it is sensitive to environmental changes, such as the movement of people.
The dataset includes seven distinct scenarios (S1-S7):
  • S1–S3: A bedroom with a central bookcase. S1 was used for model training, S2 for a generalization scenario, and S3 was measured on different days than S1;
  • S4–S5: Activities performed alternately within the bedroom;
  • S6: Data collected in a living room on different days;
  • S7: A laboratory with numerous desks, computers, and monitors, measured over several days.
Scenario S7 was chosen as the benchmark for this study’s model evaluation because it presented the greatest challenge. According to the SHARP authors, the high density of obstacles and reflective surfaces in this environment significantly impacts CSI signal quality and classification performance.
III-B Decision-Making Process and Algorithm Steps
To understand the purpose of our work, it is essential to first comprehend the original SHARP method. The SHARP process begins with the collection of CSI data from three individuals, with each of the eight activities sampled for 120 seconds. After the CSI data is extracted from the Wi-Fi 2.4 GHz packet preambles, it undergoes three key stages, as shown in Figure  1:
  1. 1.
    Pre-processing: This stage sanitizes the CSI data by eliminating noise, interference, and phase ambiguity. A reference signal path is used to remove phase offsets, recovering a clean signal that reveals the Doppler traces associated with movement;
  2. 2.
    Classification: The sanitized Doppler traces are fed into an Inception neural network. To improve generalization beyond the original architecture’s spatial feature learning, the network was enhanced with a BiLSTM layer, creating a hybrid Inception-BiLSTM model. This modification allows the model to capture both spatial and temporal dependencies, significantly boosting performance in human activity recognition;
  3. 3.
    Post-processing: The original SHARP method used a majority voting mechanism among the four receiving antennas to determine the final activity label. This step smoothed outliers and improved robustness.
In this work, we modified the post-processing phase by introducing a Support Vector Machine (SVM) before the antenna fusion stage. This ensemble learning strategy leverages the SVM to enhance classification boundaries, thereby improving the overall precision and reliability of the final decision.
Figure 1: The sequential process diagram illustrates the IBIS steps: collection of raw CSI data, passing through sanitization, neural network training, and Ensemble Learning post-Processing.
III-C Hybrid Inception-BiLSTM (IBIS) Network
To interpret and classify data from Doppler traces, we designed a hybrid neural architecture combining two networks. The Inception network processes global features and learns spatial characteristics, while the BiLSTM network captures temporal dependencies and patterns. This combination is designed to optimize both classification performance and generalization ability.
Figure 2 illustrates the integrated architecture of these two neural structures. While some layers like Normalization, Reshape, and Dropout are omitted for visualization, they are described in the text. Table II provides the input and output dimensions and parameter counts for each block. The shaded and hatched purple areas in the figure highlight the novel components of our proposal. Our approach enhances the standard Inception network, which typically only includes activation and convolution layers for learning local features. We added a MaxPooling layer and Batch Normalization to improve convergence and filter relevant features. Furthermore, a BiLSTM network was incorporated to enable the model to learn temporal patterns and motion dynamics from Doppler shifts.
III-C1 Inception Network
The Inception network is configured to extract spatial features from the Doppler trace data. A description of each block and its function is as follows:
  • Input: The entry point for the Doppler trace data;
  • Normalization: Normalizes input data by subtracting the mean and dividing by the standard deviation to stabilize training and improve convergence speed;
  • Conv2D (Conv Block): This block combines Conv2D, Batch Normalization, and an Activation function. The Conv2D layer uses filters to extract local patterns like edges and textures;
  • Batch Normalization: Normalizes activations to accelerate training and regularize the model;
  • Activation: Introduces non-linearity, allowing the model to learn complex representations;
  • Max Pooling: Reduces spatial resolution and complexity to focus on the most relevant features;
  • Concatenate: Merges activation maps from different paths, maintaining important information while reducing dimensionality;
  • Dropout: Randomly deactivates neurons during training to reduce overfitting and improve generalization;
  • Reshape: Transforms the output tensor into a sequence format, preparing it for temporal processing by the BiLSTM layer.
The combination of Conv2D, Batch Normalization, and Activation is repeated nine times within the Inception module, a core characteristic that increases depth and improves generalization by learning more abstract features.
III-C2 BiLSTM Network
The second part of the architecture is the BiLSTM network, which was included to capture crucial temporal patterns and dependencies from the Doppler traces. These traces are time-series signals that reflect movement and channel variations, often exhibiting complex patterns that the network must recognize.
The BiLSTM network’s components are described as follows:
  • BiLSTM + Attention: This layer processes input sequences in both forward and backward directions, capturing complex temporal relationships. The attention mechanism enhances the model’s focus on the most relevant time steps;
  • Global Average Pooling: Reduces the dimensionality of the feature space in preparation for the final classification layer;
  • Dense Layer: A fully connected layer that generates the final output, with the number of units depending on the number of output classes (Table II).
While the CNN-ABLSTM model from a comparative study [8] has a lower computational cost due to its lighter architecture, it only achieves an accuracy of 85.82%, compared to the 98.95% of our IBIS method. This performance gap is attributed to the limited number of attention layers in their model, which are insufficient to capture the complex temporal and spatial patterns and subtle variations present in data with high variability and Doppler effects. Our hybrid Inception-BiLSTM network, with its greater depth and robust attention mechanisms, is better equipped to handle complex local features, leading to higher learning capacity and a more efficient convergence rate. This is because the Inception module learns spatial features, while the BiLSTM module captures temporal dependencies, making the combined model more robust to noise and interference, which is ideal for classifying Doppler traces.
Figure 2: Architecture of the IBIS hybrid neural network for Human Activity Recognition using Doppler traces, showing all convolutional and attention layers.
TABLE II: Inception-BiLSTM Architecture (Compact Notation)
Block / Layer Output Shape Parameters
Input Layer (32, 32, 3) -
Normalization (32, 32, 3) mean, variance
Conv2D(3×1×3) (32, 32, 3) 3
BatchNormalization (32, 32, 3) 12
Activation (swish) (32, 32, 3) -
Conv2D(6×2×3) (31, 31, 6) 78
BatchNormalization (31, 31, 6) 24
Activation (swish) (31, 31, 6) -
Conv2D(5×2×6) (30, 30, 5) 125
BatchNormalization (30, 30, 5) 20
Conv2D(9×4×6) (28, 28, 9) 873
BatchNormalization (28, 28, 9) 36
Activation (swish) (30×30×5) / (28×28×9) -
MaxPooling2D(2×2) (15, 15, 5) -
Concatenate (15, 15, 14) -
Conv2D(3×1×14) (15, 15, 3) 45
BatchNormalization (15, 15, 3) 12
Activation (swish) (15, 15, 3) -
Conv2D(6×2×3) (14, 14, 6) 78
BatchNormalization (14, 14, 6) 24
Conv2D(5×2×6) (13, 13, 5) 125
BatchNormalization (13, 13, 5) 20
Conv2D(9×4×6) (11, 11, 9) 873
BatchNormalization (11, 11, 9) 36
Activation (swish) (13×13×5) / (11×11×9) -
MaxPooling2D(2×2) (6, 6, 5) -
Concatenate (6, 6, 14) -
Dropout(0.5) (6, 6, 14) -
Conv2D(6×1×14) (6, 6, 6) 84
BatchNormalization (6, 6, 6) 24
Activation (swish) (6, 6, 6) -
Reshape (36, 6) -
Bidirectional LSTM (64 units) (36, 128) 45568
Activation (tanh) (36, 128) -
Attention (36, 128) 16640
GlobalAvgPooling1D (RNN) (128,) -
GlobalAvgPooling1D (Attention) (128,) -
Concatenate (256,) -
Dropout(0.5) (256,) -
Dense (5 classes) (5,) 1285
Note: ‘Conv2D(out×kernel×in)‘ notation refers to output channels, filter size (square or rectangular kernel), and input channels, respectively. Shapes may vary slightly depending on padding/stride.
III-D Support Vector Machine with Radial Basis Function Kernel
The hybrid IBIS network is effective at detecting and classifying both global patterns and temporal sequences. However, for accurate classification of CSI-derived data, this architecture alone is insufficient. To address this, a SVM was integrated into the post-processing stage as an ensemble learning strategy. Unlike common approaches where SVM is the primary classifier, this work uses it to refine the final classification results and improve generalization. The SVM’s ability to handle outliers and its use of decision margins contribute to better model generalization, which is particularly useful for Doppler trace data that is prone to noise and distortions. This allows the SVM to refine the final classification results, especially for samples that might be misclassified by the neural network.
A key advantage of SVM is its kernel function flexibility. The Radial Basis Function (RBF) kernel [13] [14] was chosen for this study’s Doppler data due to its ability to adapt decision boundaries to non-linear distributions. The penalty parameter C was set to 1, providing an optimal balance between training accuracy and generalization. Additionally, class balancing was applied to the SVM configuration to ensure that each class was given equal weight, which prevents the model from being biased toward the majority class and leads to a fairer, more robust classification. Figure 3 illustrates the SVM model’s application, showing five distinct classes separated by well-defined boundaries. Proper class segmentation is essential for understanding the unique features of each class, leading to improved classification accuracy.
Figure 3: Representation of the SVM algorithm applied to five movements with non-linear Data, applying the PCA method for dimensionality reduction.
IV Performance Evaluation
This section evaluates the hybrid IBIS neural network using generalization data from Scenario S7. This scenario, a laboratory room filled with desks, chairs, and monitors, was chosen as a benchmark due to its challenging environment. The numerous obstacles in this setup introduce complex signal distortions like multipath propagation, reflection, and scattering, which make the task of activity recognition particularly difficult for any classifier. We will analyze the model’s performance through its learning curves, class-wise confusion matrices, and the computational cost of the hybrid architecture.
IV-A Confusion Matrix for 5 Movements
Figure 4 compares the confusion matrices for five movements, showing the results from the Inception-only network against the proposed hybrid scheme. As seen in Figure 4a, the Inception network struggled with precision, particularly for the walking movement, which was frequently confused with running, achieving only 71% accuracy. Jumping was also often misclassified as running, with an accuracy of just 67%. This lack of precision is attributed to the similarity of Doppler patterns between these movements and the fact that the Inception network was not designed to handle temporal dependencies crucial for accurate CSI data classification.
Figure 4: Comparative analysis of the confusion matrix for five movements: original Inception versus the hybrid IBIS with SVM post-processing.
Conversely, Figure 4b shows the results after implementing our hybrid IBIS model and the SVM algorithm, which led to significantly better performance with only minor misclassifications. The improvement is due to the BiLSTM component’s ability to extract temporal patterns from the Doppler data. Furthermore, the SVM, with its non-linear kernel, proper C parameter tuning, and class balancing, refined the model’s classifications, leading to enhanced overall performance. Both confusion matrices presented refer to the post-antenna fusion phase, where classification decisions are combined via majority voting.
IV-B ROC Curve for 5 Movements
Figure 5 displays the learning ROC curve for all five classes, with a dotted diagonal line indicating the classification threshold. This curve was generated using generalization test data without antenna fusion, allowing for a direct evaluation of the algorithm’s generalization capability.
Figure 5: ROC curve analysis for five movements: comparing the original Inception with the hybrid Neural Network IBIS with SVM Post-Processing.
As shown in Figure 5a, the original model performed poorly on the walking, running, and jumping classes, which showed significant fluctuations in accuracy. In contrast, Figure 5b demonstrates that with our proposed approach, accuracy values improved and stabilized with no apparent oscillations, even without antenna fusion. This highlights the effectiveness of applying ensemble learning in this challenging scenario.
IV-C Confusion Matrix for 8 Movements
Figure 6 shows the confusion matrices for eight movements, including: standing, stand up, and arm gym. In Figure 6a, the original algorithm’s performance worsened compared to the five-movement scenario, with a higher misclassification rate for gestures like arm gym, jumping, and sitting. This is likely due to the SHARP algorithm’s sanitization process, where the probability of error increases proportionally with the number of classified gestures and bandwidth [12]. This behavior indicates that the original model struggled with overfitting and failed to generalize when more movements were added.
Figure 6: Comparative analysis of the confusion matrix for eight movements: original Inception versus the hybrid IBIS with SVM post-processing.
In contrast, Figure 6b demonstrates that the overall accuracy remained high despite the addition of new movements. This superior performance is attributed to the combination of the IBIS network and the SVM algorithm. Even after attempting to improve the original model’s performance by modifying the dataset split to 80% for training, 10% for testing, and 10% for validation, the results remained unchanged.
IV-D ROC Curve for 8 Movements
Figure 7 presents the ROC curves for all classes in the eight-movement scenario. The original simulation, shown in Figure 7a, performed poorly on the running and jumping classes, which exhibited constant oscillations and lower accuracy. This behavior is attributed to the model’s inability to generalize effectively, resulting in misclassifications and a higher incidence of false positives.
Figure 7: ROC curve analysis for eight movements: comparing the original Inception with the hybrid IBIS approach with SVM Post-Processing.
On the other hand, Figure 7b demonstrates significantly improved results with our proposed method. The curves show no apparent oscillations and maximum accuracy was achieved across all classes. These results confirm that combining the Inception-BiLSTM architecture with the SVM algorithm is highly effective, as the performance was considerably enhanced even without antenna fusion.
IV-E Comparative Algorithms
Table III demonstrates that the hybrid Inception-BiLSTM + SVM model achieved the best and most consistent results compared to other ensemble learning models. Its superior performance is attributed to the neural network’s ability to capture complex patterns and the SVM’s capacity to handle temporal sequence boundaries, both of which are essential for Doppler-based data. The original Inception model, for example, struggled with classifying similar movements like walking and jumping, leading to false positives and poor generalization.
TABLE III: Models Accuracy
Models Empty(%) Sitting(%) Walking(%) Running(%) Jumping(%)
IBIS+SVM 100 100 98.69 97.41 97.53
Inception 100 100 70.79 100 66.67
Note: SVM - Support Vector Machine
Similarly, Table IV, which presents average metrics for five movement classes, confirms that our IBIS + SVM model outperformed other tested models. We included additional metrics like F1-score, recall, and precision to provide a more comprehensive evaluation, as relying solely on accuracy can be misleading. These metrics help assess the model’s performance in imbalanced scenarios and its ability to balance precision and recall.
TABLE IV: Average Models Metrics
Models Accuracy(%) F1 Score(%) Recall(%) Precision(%)
IBIS+SVM 98.95 98.78 98.72 98.85
Inception 88.12 88.01 87.49 91.52
Note: SVM - Support Vector Machine
Our method also showed a significant improvement over a transfer learning approach using a hybrid CNN-ABLSTM model from a related study[8]. Our proposed IBIS + SVM achieved an accuracy of 98.95% compared to their 85.82%, a remarkable 18.62% improvement under the same condition of 100 training epochs. This is because our approach uses a combination of signal processing (the SHARP algorithm for noise elimination), deep learning (the enhanced hybrid network for feature and temporal learning), and traditional machine learning (the SVM for refining classification). This pipeline and the inclusion of more attention layers, increased architectural complexity, and Doppler traces features, significantly enhanced generalization compared to the other study, which used raw unprocessed data.
IV-F Computational Complexity
A primary challenge with neural networks is computational cost, as a deeper network with more layers typically consumes more resources. To objectively measure this, Figure 8 shows the computational cost over 100 training epochs without using callbacks like early stopping.
Figure 8: Graph illustrating the execution time per training epoch for both the original Inception architecture and the proposed hybrid IBIS model.
We evaluated two distinct architectures:
  • The original Inception network;
  • An Inception network combined with a Bi-directional Long Short-Term Memory (BiLSTM) layer.
The standalone Inception network had the lowest computational cost and the shortest training time per epoch due to its shallow architecture. However, it failed to accurately classify movements, producing false positives and showing temporal instability.
V Conclusion
This study successfully addresses the challenge of Human Activity Recognition (HAR) using Wi-Fi sensing, a non-intrusive and low-cost technology with broad applications in healthcare, surveillance, and monitoring. We introduced a novel hybrid neural network architecture, IBIS, that combines Inception-BiLSTM with a SVM to enhance data generalization and classification accuracy. Our model’s performance was rigorously evaluated across scenarios with five and eight activities, achieving an exceptional average accuracy of nearly 99%. This not only demonstrates the method’s robustness against noise and interference but also represents a significant improvement of approximately 10.94% over the state-of-the-art SHARP study. The excellent results were further validated by high-performing ROC curves and other metrics.
We also conducted a comparative analysis, confirming that the Inception-BiLSTM + SVM combination outperformed other configurations. This underscores the effectiveness of our approach, which strategically leverages convolutional layers for feature extraction, BiLSTM for temporal analysis, and SVM for refining decision boundaries. For future work, we propose exploring refinements in the Doppler curves acquisition process, so that the complexity of the neural network training/post-processing can be reduced.

### v2

@@
-IBIS: A Powerful Hybrid Architecture for Human Activity Recognition
+IBIS: A Hybrid Inception-BiLSTM and SVM Ensemble for Robust Doppler-based Human Activity Recognition
 Abstract
-The increasing interest in Wi-Fi sensing stems from its potential to capture environmental data in a low-cost, non-intrusive way, making it ideal for applications like healthcare, space occupancy analysis, and gesture-based IoT control.
-However, a major limitation in this field is the common problem of overfitting, where models perform well on training data but fail to generalize to new data.
-To overcome this, we introduce a novel hybrid architecture that integrates Inception-BiLSTM with a Support Vector Machine (SVM), which we refer to as IBIS.
-Our IBIS approach is uniquely engineered to improve model generalization and create more robust classification boundaries.
-By applying this method to Doppler-derived data, we achieve a movement recognition accuracy of nearly 99%.
-Comprehensive performance metrics and confusion matrices confirm the significant effectiveness of our proposed solution.
-I Introduction
+Wi-Fi sensing is a leading technology for Human Activity Recognition (HAR), offering a non-intrusive and cost-effective solution for healthcare and smart environments.
+Despite its potential, existing methods struggle with domain shift issues, often failing to generalize to unseen environments due to overfitting.
+This paper proposes IBIS, a robust ensemble framework combining Inception-Bidirectional Long Short-Term Memory (BiLSTM) for feature extraction and Support Vector Machine (SVM) for classification of Doppler signatures.
+The proposed architecture specifically targets generalization capabilities.
+Experimental results on multiple datasets show that IBIS achieves 95.40% accuracy, delivering a 7.58% performance gain compared to standard architectures in cross-scenario evaluations on external datasets.
+The analysis confirms that IBIS effectively mitigates environmental dependency in Wi-Fi-based HAR.
+{IEEEkeywords} IBIS, Channel State Information, Human Activity Recognition, Support Vector Machine, Inception, BiLSTM, Doppler traces, Ensemble Learning.
+1 Introduction
 Wi-Fi sensing is an increasingly important technology that leverages ubiquitous Wi-Fi signals for non-intrusive Human Activity Recognition (HAR) [1].
@@
 Furthermore, raw CSI data frequently contains noise from hardware artifacts, such as phase and frequency oscillations, which can significantly hinder accurate movement interpretation and requires a rigorous sanitization process.
-To address these limitations, we propose a novel hybrid neural network architecture named Inception-BiLSTM (IBIS).
-The Inception component efficiently extracts local features from the input data [5], while the Bidirectional Long Short-Term Memory (BiLSTM) component captures long-range temporal dependencies.
-To further enhance the model’s predictive power and refine classification boundaries, we integrate a SVM with a non-linear kernel in a post-processing stage.
-This ensemble learning approach [6] leverages the strengths of multiple algorithms to achieve a more robust and reliable solution.
-Our model’s effectiveness was evaluated on five and eight human activity classification tasks using Doppler-derived data.
-We compared our IBIS approach against two state-of-the-art methods: the SHARP algorithm [7], a recent ensemble-based method for HAR, and an Attention-Based BiLSTM (ABLSTM) approach [8] that uses raw CSI.
-The results demonstrate that our hybrid Inception-BiLSTM network, augmented with a non-linear SVM, significantly outperforms both benchmark models.
-Our method achieved an average accuracy of nearly 99%, validating the benefits of combining Doppler features with a sophisticated hybrid deep learning architecture for enhanced HAR performance, particularly in challenging generalization scenarios.
-The remainder of this paper is structured as follows.
-Section II presents a comprehensive overview of existing research on HAR using hybrid neural networks, providing a fresh perspective on the field.
-Section III details the development and implementation of our proposed algorithms, including the specific neural network architecture.
-The validation process, which includes activity classification confusion matrices and Receiver Operating Characteristic (ROC) curve analysis, is described in Section IV.
-Finally, Section V provides a summary of our findings and offers concluding remarks.
-II Related Works
-This section provides a review of state-of-the-art research in Wi-Fi sensing, focusing on the machine learning algorithms commonly used for CSI data analysis.
-We also discuss ensemble learning techniques that have been applied to optimize model performance.
 Recent studies have explored various approaches to enhance Wi-Fi sensing capabilities.
-Table I presents a comparative overview of the models and key features from previous studies on CSI-based Human Activity Recognition.
 Son et al.
-[9] improved indoor occupancy detection by proposing a preprocessing method that uses nonlinear transformations to expand CSI data into three channel vectors, which amplified signal differences.
+[5] improved indoor occupancy detection by proposing a pre-processing method that uses nonlinear transformations to expand CSI data into three channel vectors, which amplified signal differences.
 A CNN was then applied, achieving an average accuracy of approximately 95%.
-This work highlights the importance of effective signal preprocessing for boosting classification performance.
+This work highlights the importance of effective signal pre-processing for boosting classification performance.
 Ensemble learning has emerged as a key strategy to improve the reliability and accuracy of machine learning models.
@@
 [6] used an ensemble learning approach for HAR by combining feature extraction with an SVM classifier.
-This method reportedly outperformed other literature approaches, reaching an accuracy of about 98.95%.
 Similarly, Ye et al.
@@
 However, they reported difficulties in reproducing the SHARP algorithm’s results, suggesting a sensitivity to environmental variability.
-Their findings also indicated that SHARP’s performance degrades with increased bandwidth and a higher number of activity classes.
-Despite these advancements, a major challenge in Wi-Fi sensing remains the issue of poor generalization to unseen data.
-He et al.
-[8] addressed this limitation by introducing a cross-subject transfer learning method based on a CNN-ABLSTM model.
-By training on separate source and target domains, they achieved an accuracy of 85.82% in cross-subject recognition tasks, demonstrating a promising path for creating more robust and generalizable models.
-TABLE I: Comparison of Models and Features in Human Activity Recognition Using CSI
-Specification [9] [10] [7] [12] [11] [8]
-Model CNN CNN+RF+SVM Inception Inception DNN-ABLSTM CNN-ABLSTM
-Accuracy 95% 98.85% 88.12% 77.72% 92.06% 85.82%
-Application Indoor Occupancy HAR HAR HAR HAR HAR
-Network 802.11 802.11n 802.11ac 802.11ac/ax 802.11n -
-Bandwidth 20MHz - 80MHz 80/160MHz - -
-CSI Data Amplitude Temporal and Spatial Correlation Phase CFR Phase CFR Amplitude and Phase Raw CSI
-III System Modeling
-III-A Dataset
-The dataset for this research is based on the SHARP study [7] and consists of CSI data collected from three volunteers (one male, two females) performing eight activities in various indoor environments.
-These activities: empty, sitting, stand up, walking, running, jumping, and arm gym—were each recorded for 120 seconds.
+Their findings also indicated that SHARP’s performance degrades with decreased bandwidth and a higher number of activity classes.
+Despite recent advances, Wi-Fi sensing still faces limited generalization to unseen data.
+Prior work, such as the CNN-ABLSTM cross-subject transfer learning method by He et al.
+[10], achieved 85.82% accuracy but highlighted the need for more robust models.
+Moreover, most existing studies rely only on the CSI amplitude, overlooking phase information, which is highly sensitive to noise and interference.
+To address these gaps, we propose IBIS, a hybrid Inception-BiLSTM architecture combined with a non-linear SVM in the post-processing stage.
+Unlike conventional ensemble approaches that fuse models during training, IBIS leverages the probability outputs of internal neural layers as high-level feature vectors for the SVM, enabling sharper decision boundaries.
+Our framework operates solely on Doppler traces extracted from the CSI phase component, exploiting information typically ignored in the literature.
+Expanding the introductory analysis presented in [11], IBIS was evaluated on five- and eight-class activity recognition tasks and compared against two state-of-the-art baselines: the SHARP algorithm [8] and an ABLSTM model using raw CSI [10].
+To assess robustness across environments, we also included a complementary Doppler-based dataset [9].
+The proposed method achieved an average accuracy of 95.40%, significantly outperforming all benchmark models and demonstrating strong generalization capabilities across different datasets and noise conditions.
+The main contribution of this proposal is to improve the generalization capability of CSI data with respect to HAR by implementing a combination of spatial features and temporal sequences through the application of the proposed framework.
+This approach enables higher activity recognition accuracy.
+In addition, the inclusion of a SVM algorithm in the post-processing stage further enhances the framework, thereby demonstrating its effectiveness compared to traditional deep learning–based methods.
+Table 1: Survey of Models and Features in Using CSI
+Method / Authors Year Task Model Bandwidth Network Accuracy CSI Data
+Son et al. [5] 2024 Indoor Ocupancy CNN 20 MHz 802.11 95% Amplitude
+WiAres [6] 2020 HAR CNN+RF+SVM - 802.11n 98.85% Temporal / Spatial Correlation
+SHARP [8] 2023 HAR Inception 20/40/80 MHz 802.11 88.12% Phase CFR
+Cominelli et al. [9] 2023 HAR Inception 20/40/80/160 MHz 802.11ac/ax 77.12% Phase CFR
+Ye et al. [7] 2024 HAR DNN-ABLSTM - 802.11n 92.06% Amplitude and Phase
+He et al. [10] 2025 HAR CNN-ABLSTM - - 85.82% Raw CSI
+Widar [12] 2022 Gesture Recognition CNN+RNN - - 92.7% Amplitude
+EfficientFi [13] 2022 HAR CNN 40 MHz 802.11ax 98% Amplitude
+WiGr [14] 2022 Gesture Recognition CNN+LSTM 40 MHz 802.11ax 86.8%–92.7% Amplitude
+Wi-SensiNet [15] 2024 HAR CNN+BiLSTM 20 MHz 802.11 99% Amplitude
+IBIS* 2025 HAR Inception-BiLSTM+SVM 80/160 MHz 802.11ac/ax 95.40% Phase CFR
+Note: The symbol * represents our proposal Inception - BILSTM + SVM architecture
+It is important to highlight that our proposal achieved better results compared to the related works presented in Table 1.
+In terms of accuracy, the IBIS framework outperformed the works that use only amplitude.
+The main point is that when using phase information, it must be carefully handled to avoid possible classification errors caused by interference and existing offsets.
+The remainder of this paper is organized as follows.
+Section 2 presents the definition of CSI as well as a detailed description of the datasets used in this study.
+Section 3 describes the development and implementation of the proposed algorithms, including the specific neural network architecture.
+The validation process, encompassing activity classification confusion matrices and Receiver Operating Characteristic (ROC) curve analysis, is presented in Section 4.
+Finally, Section 5 summarizes the main findings and provides concluding remarks.
+2 Channel State Information Data Description
+This section presents the definition of CSI as adopted in IEEE 802.11 standard networks and also describes the datasets used in this proposal for data collection.
+2.1 Channel State Information
+CSI is a data collection system that describes how a wireless signal propagates from the transmitter to the receiver.
+CSI captures the signal propagation characteristics in physical environments, including the combined effects of diffraction, reflection, and scattering [5].
+Adopting the IEEE 802.11 standard, it employs advanced techniques such as Multiple-Input Multiple-Output (MIMO) and Orthogonal Frequency Division Multiplexing (OFDM) at the physical layer.
+These technologies aim to improve the supported data capacity and enhance channel orthogonality in environments affected by multipath propagation, which occurs when transmitted signals between the transmitter and receiver follow different paths due to phenomena like collisions and external interference.
+For each transmitter/receiver antenna pair, CSI provides detailed information about the phase and amplitude attenuation experienced by the signal through multiple paths for each subcarrier.
+Specifically, the CSI for each subcarrier $k$ between a transmitter with $N_{t}$ antennas and a receiver with $N_{r}$ antennas is represented by the matrix $\mathbf{H}(k)\in\mathbb{C}$ [5]:
+$\mathbf{H}(k)=\begin{bmatrix}h_{11}(k)&\cdots&h_{1N_{t}}(k)\\ h_{21}(k)&\cdots&h_{2N_{t}}(k)\\ \vdots&\ddots&\vdots\\ h_{N_{r}1}(k)&\cdots&h_{N_{r}N_{t}}(k),\end{bmatrix},k\in[1,K]$ (1)
+where $K$ represents the number of subcarriers used, and $h_{rc}(k)$ denotes the complex channel gain from the $c{\text{-th}}$ transmitter antenna to the $r{\text{-th}}$ receiver antenna.
+Each element $h_{rc}(K)$ in the CSI matrix $\mathbf{H}(k)$ can be decomposed into amplitude and phase components, as described by the following:
+$h_{rc}(k)=|h_{rc}(k)|\cdot\exp\left(j\theta_{rc}(k)\right),$ (2)
+where $\left|h_{rc}(k)\right|$ is defined as the amplitude component, and $\theta_{rc}(k)$ is the phase component of $h_{rc}(k)$.
+Since CSI contains both amplitude and phase information for each subcarrier and for each pair of transmit and receive antennas, it provides high-resolution data compared to simple measurements such as RSS (Received Signal Strength).
+In this way, CSI data can be considered as a form of ”imaging” of the signal propagation environment.
+Changes in the environment, such as the presence or movement of people, directly affect Wi-Fi signal propagation.
+These alterations in the signal path also lead to variations in CSI.
+Since different people create distinct patterns, it is possible to estimate the number of individuals present by learning these patterns—this is the key concept behind people counting systems based on CSI-extracted data [5].
+2.2 Dataset A
+One of the datasets used for this research is based on the SHARP study [8] and consists of CSI data collected from three volunteers (one male, two females) performing eight activities in various indoor environments.
+These activities: empty, sitting, stand up, walking, running, jumping, and arm gym, were each recorded for 120 seconds.
 The data was gathered in a bedroom, living room, and university laboratory during two periods: April–December 2020 and January 2022.
@@
 According to the SHARP authors, the high density of obstacles and reflective surfaces in this environment significantly impacts CSI signal quality and classification performance.
-III-B Decision-Making Process and Algorithm Steps
-To understand the purpose of our work, it is essential to first comprehend the original SHARP method.
-The SHARP process begins with the collection of CSI data from three individuals, with each of the eight activities sampled for 120 seconds.
-After the CSI data is extracted from the Wi-Fi 2.4 GHz packet preambles, it undergoes three key stages, as shown in Figure  1:
-  1. 1.
-    Pre-processing: This stage sanitizes the CSI data by eliminating noise, interference, and phase ambiguity.
-A reference signal path is used to remove phase offsets, recovering a clean signal that reveals the Doppler traces associated with movement;
-  2. 2.
-    Classification: The sanitized Doppler traces are fed into an Inception neural network.
-To improve generalization beyond the original architecture’s spatial feature learning, the network was enhanced with a BiLSTM layer, creating a hybrid Inception-BiLSTM model.
-This modification allows the model to capture both spatial and temporal dependencies, significantly boosting performance in human activity recognition;
-  3. 3.
-    Post-processing: The original SHARP method used a majority voting mechanism among the four receiving antennas to determine the final activity label.
-This step smoothed outliers and improved robustness.
-In this work, we modified the post-processing phase by introducing a Support Vector Machine (SVM) before the antenna fusion stage.
-This ensemble learning strategy leverages the SVM to enhance classification boundaries, thereby improving the overall precision and reliability of the final decision.
-Figure 1: The sequential process diagram illustrates the IBIS steps: collection of raw CSI data, passing through sanitization, neural network training, and Ensemble Learning post-Processing.
-III-C Hybrid Inception-BiLSTM (IBIS) Network
-To interpret and classify data from Doppler traces, we designed a hybrid neural architecture combining two networks.
-The Inception network processes global features and learns spatial characteristics, while the BiLSTM network captures temporal dependencies and patterns.
-This combination is designed to optimize both classification performance and generalization ability.
+2.3 Dataset B
+For application in the experiment, we considered using a second dataset, developed by Cominelli et al.
+[9].
+Although originally designed for IEEE 802.11ax (Wi-Fi 6), thereby expanding the application scope.
+The main characteristics of this dataset are summarized below.
+This dataset was developed using four Asus RT-AX86U routers interconnected to form a local-area network (LAN) and controlled by a host notebook.
+One of the routers is configured to generate dummy Wi-Fi traffic at a constant rate through feature injection for AX-CSI, while the other three extract the corresponding CSI.
+According to the authors, due to the use of IEEE 802.11ax, data grows very rapidly, making it necessary to store the recordings in PCAP files.
+Data collection was performed across seven different scenarios:
+  • S1 to S5: Conducted in the same physical environment, where neither the receivers nor the transmitter were moved during the entire duration of the experiments;
+  • S6: An office environment containing chairs and desks;
+  • S7: A large hall, representing an open environment without obstacles.
+Twelve distinct activities were recorded: Walk, Run, Jump, Sitting, Empty room, Standing, Wave hands, Clapping, Lay down, Wiping, Squat, Stretching with a sampling rate of 80 samples per second.
+For our experiment, scenario S2 was selected. This environment consists of a laboratory with reflective surfaces, which makes prediction more difficult.
+Due to the accuracy challenges encountered with the original model, our intention is to apply the IBIS framework in different environments and scenarios in order to evaluate its behavior and robustness.
+3 System Modeling
+In this section, the main definitions related to the IBIS framework are presented, including the architectural composition, as well as the derived pseudocode outlining the main steps of the framework operation.
+3.1 Decision-Making Process and Algorithm Steps
+To understand the purpose of our work, it is essential to first comprehend the SHARP method.
+After the CSI data is extracted from the Wi-Fi 802.11 ac/ax packet preambles, it undergoes three key stages, as shown in Figure  1:
+3.1.1 Pre-processing
+This stage sanitizes the CSI data by eliminating noise, interference, and phase ambiguity.
+A reference signal path is used to remove phase offsets, recovering a clean signal that reveals the Doppler traces associated with movement.
+3.1.2 Classification
+The sanitized Doppler traces are fed into an Inception neural network.
+To improve generalization beyond the original architecture’s spatial feature learning capabilities, the network was enhanced with a BiLSTM layer, creating a hybrid Inception–BiLSTM model.
+This modification enables the architecture to capture both spatial and temporal dependencies, significantly boosting performance in human activity recognition.
+At this stage, it is possible to select the number of antennas used in the training process.
+The more antennas that are selected—up to the limit of four—the more distinctive information becomes available, consequently improving the neural network’s recognition performance.
+To understand how the IBIS operates, it is important to recognize the characteristics and advantages introduced by this new configuration.
+Unlike the original Inception network, additional layers are incorporated, providing enhanced data interpretation and improving the model’s ability to extract discriminative features.
 Figure 7 illustrates the integrated architecture of these two neural structures.
 While some layers like Normalization, Reshape, and Dropout are omitted for visualization, they are described in the text.
-Table II provides the input and output dimensions and parameter counts for each block.
+Table 2 provides the input and output dimensions and parameter counts for each block.
 The shaded and hatched purple areas in the figure highlight the novel components of our proposal.
@@
 Furthermore, a BiLSTM network was incorporated to enable the model to learn temporal patterns and motion dynamics from Doppler shifts.
-III-C1 Inception Network
-The Inception network is configured to extract spatial features from the Doppler trace data. A description of each block and its function is as follows:
+Inception Network: This module is configured to extract spatial features from the Doppler trace data. A description of each block and its function is as follows:
   • Input: The entry point for the Doppler trace data;
@@
 The combination of Conv2D, Batch Normalization, and Activation is repeated nine times within the Inception module, a core characteristic that increases depth and improves generalization by learning more abstract features.
-III-C2 BiLSTM Network
-The second part of the architecture is the BiLSTM network, which was included to capture crucial temporal patterns and dependencies from the Doppler traces.
+BiLSTM Network: The second part of the architecture is the BiLSTM network, which was included to capture crucial temporal patterns and dependencies from the Doppler traces.
 These traces are time-series signals that reflect movement and channel variations, often exhibiting complex patterns that the network must recognize.
@@
   • Global Average Pooling: Reduces the dimensionality of the feature space in preparation for the final classification layer;
-  • Dense Layer: A fully connected layer that generates the final output, with the number of units depending on the number of output classes (Table II).
-While the CNN-ABLSTM model from a comparative study [8] has a lower computational cost due to its lighter architecture, it only achieves an accuracy of 85.82%, compared to the 98.95% of our IBIS method.
-This performance gap is attributed to the limited number of attention layers in their model, which are insufficient to capture the complex temporal and spatial patterns and subtle variations present in data with high variability and Doppler effects.
-Our hybrid Inception-BiLSTM network, with its greater depth and robust attention mechanisms, is better equipped to handle complex local features, leading to higher learning capacity and a more efficient convergence rate.
-This is because the Inception module learns spatial features, while the BiLSTM module captures temporal dependencies, making the combined model more robust to noise and interference, which is ideal for classifying Doppler traces.
-Figure 2: Architecture of the IBIS hybrid neural network for Human Activity Recognition using Doppler traces, showing all convolutional and attention layers.
-TABLE II: Inception-BiLSTM Architecture (Compact Notation)
+  • Dense Layer: A fully connected layer that generates the final output, with the number of units depending on the number of output classes (Table 2).
+3.1.3 Post-processing
+The original SHARP method used a majority voting mechanism among the four receiving antennas to determine the final activity label. This step smoothed outliers and improved robustness.
+In this work, we modified the post-processing phase by introducing a Support Vector Machine (SVM) before the antenna fusion stage.
+This ensemble learning strategy leverages the SVM to enhance classification boundaries, thereby improving the overall precision and reliability of the final decision.
+Figure 1: The sequential process diagram illustrates the IBIS steps: collection of raw CSI data, passing through sanitization, neural network training, and Ensemble Learning post-Processing.
+Figure 2: Architecture of the Inception-BiLSTM hybrid neural network for Human Activity Recognition using Doppler traces, showing all convolutional and attention layers.
+Table 2: Inception-BiLSTM Architecture (Compact Notation)
 Block / Layer Output Shape Parameters
@@
 Note: ‘Conv2D(out×kernel×in)‘ notation refers to output channels, filter size (square or rectangular kernel), and input channels, respectively. Shapes may vary slightly depending on padding/stride.
-III-D Support Vector Machine with Radial Basis Function Kernel
-The hybrid IBIS network is effective at detecting and classifying both global patterns and temporal sequences.
+3.2 Post-Processing Support Vector Machine
+The hybrid Inception-BiLSTM neural network is effective at detecting and classifying both global patterns and temporal sequences.
 However, for accurate classification of CSI-derived data, this architecture alone is insufficient.
-To address this, a SVM was integrated into the post-processing stage as an ensemble learning strategy.
+To address this, an SVM was integrated into the post-processing stage as an ensemble learning strategy.
 Unlike common approaches where SVM is the primary classifier, this work uses it to refine the final classification results and improve generalization.
@@
 This allows the SVM to refine the final classification results, especially for samples that might be misclassified by the neural network.
-A key advantage of SVM is its kernel function flexibility.
-The Radial Basis Function (RBF) kernel [13] [14] was chosen for this study’s Doppler data due to its ability to adapt decision boundaries to non-linear distributions.
-The penalty parameter C was set to 1, providing an optimal balance between training accuracy and generalization.
-Additionally, class balancing was applied to the SVM configuration to ensure that each class was given equal weight, which prevents the model from being biased toward the majority class and leads to a fairer, more robust classification.
+A key advantage of SVM lies in the flexibility of its kernel functions.
+The Radial Basis Function (RBF) kernel, as well as the Sigmoid and Polynomial kernels [16][17][18], are available options that can be selected for classification.
+In our framework, during the training stage, the SVM is trained using the probability outputs generated by the neural network.
+Through a grid-search optimization process, the most suitable kernel is determined for the underlying class distribution.
+The penalty parameter C is also tuned to achieve an optimal balance between training accuracy and model generalization.
+Furthermore, class balancing was applied to the SVM configuration to ensure that all classes receive equal weighting, preventing the classifier from becoming biased toward majority classes and resulting in a fairer and more robust overall prediction performance.
 Figure 3 illustrates the SVM model’s application, showing five distinct classes separated by well-defined boundaries.
 Proper class segmentation is essential for understanding the unique features of each class, leading to improved classification accuracy.
+Regarding the hyperparameters, a learning rate of 0.001 was used for the hybrid network, along with an early stopping callback function to achieve the highest possible accuracy.
+At the classification stage, the SVM was also trained using a Grid Search, where three parameters were varied to determine the optimal training configuration.
+Initially, three kernel functions were considered: RBF, sigmoid, and polynomial, along with the trade-off parameter $C$ and gamma ($\gamma$), both ranging from 0.01 to 1, which are responsible for defining the decision margins.
+The RBF kernel was ultimately selected by the Grid Search, representing the non-linear characteristics of the data.
+The main reason for incorporating an SVM in the post-processing stage is that the Softmax layer of a neural network tends to force a probabilistic decision, even in highly ambiguous classes, and is sensitive to small variations in previously unseen environments.
+In contrast, the SVM aims to maximize the separation margin using high-level feature vectors, receiving a cleaner representation of class boundaries.
+This makes the inclusion of an SVM particularly suitable for noisy Doppler traces, where decision boundaries are not well defined.
+In this context, the combination of the BiLSTM module with dropout regularization, followed by the SVM in the post-processing step, contributes significantly to improving the model’s generalization capability.
 Figure 3: Representation of the SVM algorithm applied to five movements with non-linear Data, applying the PCA method for dimensionality reduction.
-IV Performance Evaluation
-This section evaluates the hybrid IBIS neural network using generalization data from Scenario S7.
-This scenario, a laboratory room filled with desks, chairs, and monitors, was chosen as a benchmark due to its challenging environment.
+3.3 Pseudocode
+The pseudocode description presented in algorithm  1 concisely illustrates the application of the IBIS framework and the main processes involved.
+It is worth mentioning that the code containing the modifications to the SHARP algorithm, as well as all stages of the IBIS implementation, will be made public after the publication of this proposal.
+Algorithm 1 IBIS framework
+1:Preprocessing:
+2:    Apply LASSO-based sanitization and remove offset
+3:    Extract Doppler traces and select subcarriers
+4:Input: Dataset $X$, labels $Y$, batch size 3232, learning rate $0.001$, epochs 100100
+5:    Normalize features: $X_{\text{norm}}=\text{normalize}(X)$
+6:    Split dataset: $(X_{train},Y_{train},X_{test},Y_{test})\leftarrow\text{split}(X_{\text{norm}},Y,0.8)$
+7:Inception Feature Extraction:
+8:    Apply parallel convolutions with multiple kernel sizes
+9:    Concatenate outputs to obtain $F_{\text{inception}}$
+10:BiLSTM Temporal Modeling:
+11:    Feed $F_{\text{inception}}$ into BiLSTM layers
+12:    Extract temporal representation $F_{\text{BiLSTM}}$
+13:Softmax Classification:
+14:    Compute posterior probabilities:
+15:     $P(Y|X)=\text{Softmax}(W_{s}F_{\text{BiLSTM}}+b_{s})$
+16:    Perform Grid Search to select kernel type, $C$, and $\gamma$
+17:SVM Training (Postprocessing):
+18:    Train SVM using $(P(Y|X_{train}),Y_{train})$
+19:    Available kernels:
+20:     RBF: $K(x,x^{\prime})=\exp(-\gamma\|x-x^{\prime}\|^{2})$
+21:     Sigmoid: $K(x,x^{\prime})=\tanh(\gamma\,x^{\top}x^{\prime}+r)$
+22:     Polynomial: $K(x,x^{\prime})=(\gamma\,x^{\top}x^{\prime}+r)^{d}$
+23:Prediction:
+24:    Compute $P(Y|X_{test})$ using the Inception–BiLSTM–Softmax network
+25:    Predict: $Y_{pred}=\text{SVM.predict}(P(Y|X_{test}))$
+26:Evaluation:
+27:    Compute Accuracy, Precision, Recall, and F1-score
+28:Output: Final predictions $Y_{pred}$ and evaluation metrics
+4 Performance Evaluation
+This section evaluates the IBIS framework using generalization data from Scenario S7 for Dataset A and from Scenario S2 for Dataset B, as previously described.
+It is important to note that this validation used an 80 MHz bandwidth (IEEE 802.11ac) for Scenario S7 and a 160 MHz bandwidth (IEEE 802.11ax) for Scenario S2, with the selection of four antennas to maximize the amount of information obtained during training.
+This scenario, a laboratory room filled with desks, chairs, and monitors, was chosen as a benchmark due to its challenging nature.
 The numerous obstacles in this setup introduce complex signal distortions like multipath propagation, reflection, and scattering, which make the task of activity recognition particularly difficult for any classifier.
 We will analyze the model’s performance through its learning curves, class-wise confusion matrices, and the computational cost of the hybrid architecture.
-IV-A Confusion Matrix for 5 Movements
+For data separation, after Doppler extraction, the dataset was divided into 60%, 20%, and 20% for training, testing, and validation, respectively.
+To avoid decision bias in the evaluated algorithms, ten simulations were performed for each experiment, and the results provided herein represent the average across the ten simulations.
+4.1 Confusion Matrix for 5 Movements
 Figure 4 compares the confusion matrices for five movements, showing the results from the Inception-only network against the proposed hybrid scheme.
 As seen in Figure 4a, the Inception network struggled with precision, particularly for the walking movement, which was frequently confused with running, achieving only 71% accuracy.
-Jumping was also often misclassified as running, with an accuracy of just 67%.
+Jumping was also often misclassified as running, with an accuracy of just 70%.
 This lack of precision is attributed to the similarity of Doppler patterns between these movements and the fact that the Inception network was not designed to handle temporal dependencies crucial for accurate CSI data classification.
-Figure 4: Comparative analysis of the confusion matrix for five movements: original Inception versus the hybrid IBIS with SVM post-processing.
-Conversely, Figure 4b shows the results after implementing our hybrid IBIS model and the SVM algorithm, which led to significantly better performance with only minor misclassifications.
-The improvement is due to the BiLSTM component’s ability to extract temporal patterns from the Doppler data.
-Furthermore, the SVM, with its non-linear kernel, proper C parameter tuning, and class balancing, refined the model’s classifications, leading to enhanced overall performance.
-Both confusion matrices presented refer to the post-antenna fusion phase, where classification decisions are combined via majority voting.
-IV-B ROC Curve for 5 Movements
+(a) Inception - 88.16%
+(b) IBIS - 95.40%
+Note: E - Empty, S - Sitting, W - Walking, R - Running, and J - Jumping
+Figure 4: Comparative analysis of the confusion matrix for five movements: original Inception versus IBIS
+Conversely, Figure 4b shows the results obtained after implementing our IBIS framework, which achieved substantially better performance with only minor misclassifications.
+A more balanced distribution of accuracy across the classes is also observed, reaching up to 99% in categories where the original model previously exhibited confusion.
+Overall, this configuration yields a 7.58% performance gain compared to the previous simulation.
+This improvement is primarily attributed to the BiLSTM component’s ability to capture temporal patterns from the Doppler data.
+In addition, the SVM—through the selection of an appropriate kernel, proper tuning of the parameter $C$, and class balancing—further refined the classification boundaries, resulting in enhanced overall performance.
+Both confusion matrices shown correspond to the post-antenna fusion stage, where classification decisions are merged using majority voting.
+4.2 ROC Curve for 5 Movements
 Figure 5 displays the learning ROC curve for all five classes, with a dotted diagonal line indicating the classification threshold.
 This curve was generated using generalization test data without antenna fusion, allowing for a direct evaluation of the algorithm’s generalization capability.
-Figure 5: ROC curve analysis for five movements: comparing the original Inception with the hybrid Neural Network IBIS with SVM Post-Processing.
-As shown in Figure 5a, the original model performed poorly on the walking, running, and jumping classes, which showed significant fluctuations in accuracy.
-In contrast, Figure 5b demonstrates that with our proposed approach, accuracy values improved and stabilized with no apparent oscillations, even without antenna fusion.
-This highlights the effectiveness of applying ensemble learning in this challenging scenario.
-IV-C Confusion Matrix for 8 Movements
-Figure 6 shows the confusion matrices for eight movements, including: standing, stand up, and arm gym.
-In Figure 6a, the original algorithm’s performance worsened compared to the five-movement scenario, with a higher misclassification rate for gestures like arm gym, jumping, and sitting.
-This is likely due to the SHARP algorithm’s sanitization process, where the probability of error increases proportionally with the number of classified gestures and bandwidth [12].
+(a) Inception
+(b) IBIS
+Figure 5: ROC curve analysis for five movements: comparing the original Inception with IBIS
+As shown in Figure 5a, the original model performed poorly for the walking, running, and jumping classes, which exhibited significant fluctuations in accuracy.
+In contrast, Figure 5b shows that, with our proposed approach, the accuracy values increased and became more stable, with no apparent oscillations, even without antenna fusion, except for the Walking class.
+At this stage, we are evaluating only the effect of applying the SVM after training the neural network using the softmax probabilities as input.
+In other words, this result reflects exclusively the effectiveness of the SVM refinement, excluding the majority-voting step from antenna fusion, which is applied afterward.
+However, when considering the class distribution, the results are more uniform.
+This highlights the effectiveness of incorporating ensemble learning in this challenging scenario.
+4.3 Confusion Matrix for 8 Movements
+Figure 6 shows the confusion matrices for eight movements, including standing, stand up, and arm gym.
+In Figure 6a, the performance of the original algorithm deteriorated compared to the five-movement scenario, exhibiting a higher misclassification rate for gestures such as arm gym, jumping, and sitting, resulting in an accuracy of 38.38%.
+This is likely due to the SHARP algorithm’s sanitization process, where the error probability increases proportionally to the number of classified gestures and bandwidth [9].
 This behavior indicates that the original model struggled with overfitting and failed to generalize when more movements were added.
-Figure 6: Comparative analysis of the confusion matrix for eight movements: original Inception versus the hybrid IBIS with SVM post-processing.
-In contrast, Figure 6b demonstrates that the overall accuracy remained high despite the addition of new movements.
-This superior performance is attributed to the combination of the IBIS network and the SVM algorithm.
+(a) Inception - 38.38%
+(b) IBIS - 70.25%
+Note: E - Empty, S - Sitting, W - Walking, R - Running, J - Jumping, U - Stand Up, G - Standing and A - Arm Gym
+Figure 6: Comparative analysis of the confusion matrix for eight movements: original Inception versus IBIS
+In contrast, Figure 6b shows that the overall accuracy remained high despite the addition of new movements, reaching 70.25%.
+This superior performance is attributed to the combination of the Inception-BiLSTM hybrid network and the SVM algorithm.
 Even after attempting to improve the original model’s performance by modifying the dataset split to 80% for training, 10% for testing, and 10% for validation, the results remained unchanged.
-IV-D ROC Curve for 8 Movements
-Figure 7 presents the ROC curves for all classes in the eight-movement scenario.
-The original simulation, shown in Figure 7a, performed poorly on the running and jumping classes, which exhibited constant oscillations and lower accuracy.
-This behavior is attributed to the model’s inability to generalize effectively, resulting in misclassifications and a higher incidence of false positives.
-Figure 7: ROC curve analysis for eight movements: comparing the original Inception with the hybrid IBIS approach with SVM Post-Processing.
-On the other hand, Figure 7b demonstrates significantly improved results with our proposed method.
-The curves show no apparent oscillations and maximum accuracy was achieved across all classes.
-These results confirm that combining the Inception-BiLSTM architecture with the SVM algorithm is highly effective, as the performance was considerably enhanced even without antenna fusion.
-IV-E Comparative Algorithms
-Table III demonstrates that the hybrid Inception-BiLSTM + SVM model achieved the best and most consistent results compared to other ensemble learning models.
+4.4 Doppler Traces
+In Figure 7, we present the Doppler trace plots for eight activities.
+It is possible to observe that the traces corresponding to the Empty, Sitting, Standing, and Arm Gym activities exhibit very similar movement patterns, making them difficult for a neural network to distinguish.
+Consequently, their spatial features lie very close to one another.
+In this context, the BiLSTM network combined with the SVM becomes the decisive component: the BiLSTM is able to extract temporal dependencies, and once these dependencies are captured, the SVM refines the decision boundaries through appropriate tuning of the trade-off parameter, kernel choice, and margin.
+It becomes clear that, even with IBIS, the framework struggles to reliably distinguish highly similar Doppler traces due to the strong overlap between activities.
+This reinforces the need for a carefully designed preprocessing pipeline.
+Moreover, depending on the normalization strategy, relevant offsets and peak-related features may be suppressed, which are essential for accurate neural network inference.
+Nevertheless, IBIS still outperforms the original Inception network, yielding an improvement of approximately 45.30%.
+Figure 7: Doppler traces for the activities: E – Empty, S – Sitting, W – Walking, R – Running, J – Jumping, U – Stand Up, G – Standing, A – Arm Gym.
+4.5 Ablation
+The next experiment conducted with our framework aims to evaluate its behavior when individual components of the pipeline are removed.
+First, we run the experiment using the SVM together with the Inception network and the full IBIS model.
+In the second step, we remove the SVM from the composition, leaving only the Inception–BiLSTM neural network.
+Analyzing the first experiment in Figure 8, we observe that the Inception model with SVM experiences a reduction in performance, achieving 85.27%, whereas the full IBIS model reaches 95.40%.
+Removing the SVM from the pipeline resulted in a 2% increase in accuracy, highlighting the contribution of the BiLSTM component, which achieved 90% compared to the 88% produced by the original Inception network.
+When this performance is compared with the 95.40% obtained by the complete framework, it becomes clear that the SVM plays a predominant role, as most of the accuracy gain originates from its inclusion in the post-processing stage, where it effectively refines class boundaries with high precision.
+Another experiment involved removing the Doppler traces from the pipeline and applying IBIS directly to the sanitized data.
+Analyzing Figure 8, on the right-hand side, we observe that the Inception network without Doppler information reaches 34% accuracy, while IBIS achieves 49%.
+Although both results are lower, this confirms that Doppler features play a fundamental role in the analysis of phase components.
+For this specific dataset, removing Doppler information leads to a substantial degradation in accuracy, especially for the Inception network.
+Figure 8: Graph illustrating the behavior of the accuracy when removing the Doppler traces and/or the SVM components
+4.6 Exploring another dataset
+With the intention of demonstrating the efficiency of the IBIS framework, new experiments were conducted using the public dataset provided by Cominelli et al.
+[9].
+In this experiment, Scenario S2 was used, and—similarly to the previous experiments—the Grid Search technique was applied to select the trade-off parameter $C$ and the $\gamma$ parameter, as well as to evaluate different kernels, including RBF, Sigmoid, and Polynomial.
+Thus, the exact same procedure adopted in our framework was applied here, but now with a different dataset.
+In this previously described dataset, several SHARP adaptations were applied to support operation with a 160 MHz bandwidth.
+Five activities—Walking, Running, Jumping, Sitting, and Empty Room, were evaluated, as shown in the confusion matrix presented in Figure 9.
+Based on this dataset, adjustments were also made to the scripts to make them more consistent with those used in the other dataset.
+It is important to highlight that, unlike the previous dataset, the IBIS framework was not able to effectively distinguish between the walking and running activities in this dataset.
+This limitation is mainly due to the reduced amount of available information: after generating the Doppler traces, the window segments created during the dataset splitting and training stages become significantly smaller.
+As a result, the IBIS does not receive sufficient information to reliably differentiate between these two activities.
+Analyzing Figure 9, we observe that the Inception network alone achieved an accuracy of 70.23%, whereas the IBIS framework reached 81.40%, representing an accuracy gain of 13.72%.
+This difference in performance between datasets can be attributed to variations in data distribution; elements such as obstacles, walls, and interference introduce additional complexity and reduce recognition performance due to the increased similarity between activity patterns.
+These experiments demonstrate that IBIS maintains superior performance across different datasets, achieving substantial improvements even under varying environmental conditions and inherent limitations.
+(a) Inception - 70.23%
+(b) IBIS - 81.40%
+Note: A - Walking, B - Run, C - Jump, D - Sitting, and E - Empty Room
+Figure 9: Comparative analysis of the confusion matrices for five activities: original Inception model versus the IBIS, using the dataset from Cominelli et al. [9]
+4.7 Comparable Algorithms
+Table 3 demonstrates that the IBIS framework achieved the best and most consistent results compared to other ensemble learning models.
 Its superior performance is attributed to the neural network’s ability to capture complex patterns and the SVM’s capacity to handle temporal sequence boundaries, both of which are essential for Doppler-based data.
 The original Inception model, for example, struggled with classifying similar movements like walking and jumping, leading to false positives and poor generalization.
-TABLE III: Models Accuracy
+Table 3: Models Accuracy
 Models Empty(%) Sitting(%) Walking(%) Running(%) Jumping(%)
-IBIS+SVM 100 100 98.69 97.41 97.53
-Inception 100 100 70.79 100 66.67
-Note: SVM - Support Vector Machine
-Similarly, Table IV, which presents average metrics for five movement classes, confirms that our IBIS + SVM model outperformed other tested models.
+IBIS 100 87 91 100 99
+Inception 100 99 71 100 70
+Similarly, Table 4, which presents average metrics for five movement classes, confirms that our IBIS outperformed other tested models.
 We included additional metrics like F1-score, recall, and precision to provide a more comprehensive evaluation, as relying solely on accuracy can be misleading.
 These metrics help assess the model’s performance in imbalanced scenarios and its ability to balance precision and recall.
-TABLE IV: Average Models Metrics
+Table 4: Average Models Metrics for 5 Movements
 Models Accuracy(%) F1 Score(%) Recall(%) Precision(%)
-IBIS+SVM 98.95 98.78 98.72 98.85
-Inception 88.12 88.01 87.49 91.52
-Note: SVM - Support Vector Machine
-Our method also showed a significant improvement over a transfer learning approach using a hybrid CNN-ABLSTM model from a related study[8].
-Our proposed IBIS + SVM achieved an accuracy of 98.95% compared to their 85.82%, a remarkable 18.62% improvement under the same condition of 100 training epochs.
-This is because our approach uses a combination of signal processing (the SHARP algorithm for noise elimination), deep learning (the enhanced hybrid network for feature and temporal learning), and traditional machine learning (the SVM for refining classification).
-This pipeline and the inclusion of more attention layers, increased architectural complexity, and Doppler traces features, significantly enhanced generalization compared to the other study, which used raw unprocessed data.
-IV-F Computational Complexity
+IBIS 95.40 95.21 95.30 95.57
+Inception 88.16 88.26 87.87 91.47
+While the CNN-ABLSTM model from a related study [10] presents a lower computational cost due to its lighter architecture, it achieves only 85.82% accuracy, whereas our IBIS method reaches 95.40%.
+This 10.04% improvement under the same 100-epoch training condition is primarily due to the limited number of attention layers and the reduced architectural complexity in their model, which hinder its ability to capture complex temporal–spatial patterns and subtle Doppler-induced variations.
+In contrast, our IBIS framework leverages the Inception module for spatial feature extraction and the BiLSTM module for temporal dependency learning, making it more robust to noise, interference, and reflective-surface effects.
+Additionally, our pipeline integrates advanced signal processing, the SHARP algorithm for noise suppression, and traditional machine learning SVM for refined classification.
+This combined approach significantly improves generalization compared to the transfer-learning-based model trained on raw, unprocessed data in  [10].
+4.8 Computational Complexity
 A primary challenge with neural networks is computational cost, as a deeper network with more layers typically consumes more resources.
@@
   • The original Inception network;
-  • An Inception network combined with a Bi-directional Long Short-Term Memory (BiLSTM) layer.
+  • An Inception network combined with a BiLSTM layer.
 The standalone Inception network had the lowest computational cost and the shortest training time per epoch due to its shallow architecture.
 However, it failed to accurately classify movements, producing false positives and showing temporal instability.
-V Conclusion
-This study successfully addresses the challenge of Human Activity Recognition (HAR) using Wi-Fi sensing, a non-intrusive and low-cost technology with broad applications in healthcare, surveillance, and monitoring.
-We introduced a novel hybrid neural network architecture, IBIS, that combines Inception-BiLSTM with a SVM to enhance data generalization and classification accuracy.
-Our model’s performance was rigorously evaluated across scenarios with five and eight activities, achieving an exceptional average accuracy of nearly 99%.
-This not only demonstrates the method’s robustness against noise and interference but also represents a significant improvement of approximately 10.94% over the state-of-the-art SHARP study.
-The excellent results were further validated by high-performing ROC curves and other metrics.
-We also conducted a comparative analysis, confirming that the Inception-BiLSTM + SVM combination outperformed other configurations.
-This underscores the effectiveness of our approach, which strategically leverages convolutional layers for feature extraction, BiLSTM for temporal analysis, and SVM for refining decision boundaries.
-For future work, we propose exploring refinements in the Doppler curves acquisition process, so that the complexity of the neural network training/post-processing can be reduced.
+It should be noted that the SVM classifier is not displayed in this figure because it is not involved in the epoch-level training process.
+The SVM is fitted only once after the neural network has converged, using the extracted feature probabilities.
+Consequently, its computational overhead is negligible, and its inclusion in the graphical representation is unnecessary.
+5 Conclusion
+This paper addresses the challenges of generalization and robustness in HAR using non-intrusive Wi-Fi sensing.
+We propose IBIS, a novel hybrid architecture that strategically integrates Inception layers for spatial feature extraction, a BiLSTM module for temporal modeling, and an SVM classifier for sharper decision boundaries.
+Extensive evaluations across scenarios with different activity sets (five and eight classes) demonstrate the framework’s effectiveness, reaching an average accuracy of 95.40%.
+Notably, IBIS outperforms the state-of-the-art SHARP method by 7.58% on Dataset A, highlighting its superior capability in handling noise and environmental interference.
+A second improvement of 13.40% was observed on Dataset B, largely due to environmental differences, where variations in surface materials and layout can alter signal propagation patterns.
+Nevertheless, even under new conditions and datasets, IBIS consistently delivers superior performance.
+The analysis further confirms that replacing the standard Softmax layer with an SVM classifier significantly enhances classification performance in complex environments.
+Future work will focus on optimizing the Doppler trace acquisition and pre-processing stages to reduce computational complexity and further streamline the training pipeline.